Solve for : --x + 4x – 3 50 Type int and -inf for oo and -20, respectively. • Type the capital letter U to represent th
Answer #1

x2 + 4 - 320 4 - 30+لالا- X 4X + 3 = 0 X2-X - 3X + 3 = 6 - 3) * )2 - (ا - (3-X) ( الا) مبادلا roots 1 3